OpenStudy (anonymous):

Please help for Medal!! What is the slope of the line that passes through the points (3, –1) and (–2, –5)? A. -4/5 B.4/5 C.5/4 D.-5/4

OpenStudy (princeharryyy):

use (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Well, to find the slope of a line, you must follow this equation: \[\frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\]

OpenStudy (princeharryyy):

(-5+1)/(-2-3) = 4/5

OpenStudy (anonymous):

Plug it in: \[\frac{ -5 - (-1) }{ -2 - 3 }\]

OpenStudy (anonymous):

It'd actually be \[\frac{ -4 }{ 5 }\]

OpenStudy (princeharryyy):

-4/(-5) = 4/5 @kybi

OpenStudy (anonymous):

Yes it is 4/5 Correct @princeharryyy

OpenStudy (anonymous):

Hm, true. Sorry, I made a slight error with the negatives

OpenStudy (anonymous):

Thank you @princeharryyy :)

OpenStudy (anonymous):

Sorry for the wrong-ish answer xD

OpenStudy (princeharryyy):

no problem @kybi

OpenStudy (anonymous):

it's fine @kybi thank you for trying :)

OpenStudy (anonymous):

Assuming the points are (0,2) and (5,0) and that the "-" is used as a separator rather than as a minus sign. If it is a minus sign, why -0? Slope = difference in y coordinates/difference in x coordinates = (2 - 0)/(0 - 5) = 2/(-5) = -0.4